OpenStudy (anonymous):

If 60 feet of uniform wire weigh 80 pounds, what is the weight of 2 yards of the same wire? (A)2 2/3 (B)6. (C)2400. (D)120. (E)8

OpenStudy (gw2011):

The answer is (E) 8. Here is how I got the answer: Set up a proportion: 60/80=6/x Cross multiplying you get: 60x=(80)(6) Since there are 3 feet to 1 yard, then 2 yards = 6 feet x=8